You're shopping at Festival Foods before some friends come over for the weekend. There are 5 types of chips, 4 types of soda and 6 types of candy to pick from. How many possible combinations do you have for purchasing?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3835)


If you buy one of each category, then there would be 120 different combinations because 5*4*6 = 20*6 = 120.

Consider for a moment that you just want 1 bag of chips and 1 soda.
Make a table that has 5 rows and 4 columns. It would have 5*4 = 20 inner cells to it representing the possible paired values of chips & soda.

Now imagine taking those 20 entries to make a new table with 20 rows.
This new table will also have 6 columns.
The 20 rows represent the chips & soda pairings.
The 6 columns would be the different candy types.
This second table has 20*6 = 120 cells to represent all possible trios of chips/soda/candy.

Making such tables is optional. It could be done as a thought exercise rather than actually writing down the table.

Answer by ikleyn(53751)

.

5*4*6 = 120 different possible combinations for purchasing.

Use the Fundamental counting principle.

---------------

On Fundamental counting principle,  see the lesson
    - Fundamental counting principle problems
in this site.

Find there many other similar  (and different)  solved problems.

Read the lesson and learn the subject from there.


Happy learning  ( ! )